### v2

Abstract
Existing language models (LMs) often exhibit a Western-centric bias and struggle to represent diverse cultural knowledge. Previous attempts to address this rely on synthetic data and express cultural knowledge only in English. In this work, we study whether a small amount of human-written, multilingual cultural preference data can improve LMs across various model families and sizes. We first introduce CARE, a multilingual resource of 24.1k responses with human preferences on 2,580 questions about Chinese and Arab cultures, all carefully annotated by native speakers and offering more balanced coverage. Using CARE, we demonstrate that cultural alignment improves existing LMs beyond generic resources without compromising general capabilities. Moreover, we evaluate the cultural awareness of LMs, native speakers, and retrieved web content when queried in different languages. Our experiment reveals regional disparities among LMs, which may also be reflected in the documentation gap: native speakers often take everyday cultural commonsense and social norms for granted, while non-natives are more likely to seek out and document them actively. CARE is publicly available at https://github.com/Guochry/CARE (we plan to add Japanese data in the near future).
CARE: Aligning Language Models for Regional Cultural Awareness
[…]
1 Introduction
The impressive capabilities of multilingual language models (LMs) lead to their widespread use across various cultural groups Ouyang et al. (2022); Dubey et al. (2024). However, many studies have shown that LMs often lack knowledge of diverse cultures and are biased towards Western culture Naous et al. (2023); AlKhamissi et al. (2024); Naous and Xu (2025). While much focus has been on demonstrating these cultural biases, less effort Li et al. (2024a, b); Kirk et al. (2024) has been made to address them. Existing works have largely been limited to expressing cultural knowledge in only English or generating synthetic data with existing LMs for fine-tuning.
Figure 1: Example LM responses to culture-specific questions in Chinese and Arabic. The base LM (Refer to caption) Llama3.1-8B-Instruct fails to respond appropriately, while its aligned versions on CARE generate better responses for Chinese (Refer to caption) and Arab (Refer to caption) cultures.
In this work, we explore the benefits of learning from multilingual multicultural human preference data, and demonstrate how even a modest dataset of 2.6k high-quality human-written samples can significantly enhance LM’s cultural awareness across various model families and sizes.
Specifically, we first introduce the Cultural Alignment REsource (CARE [Uncaptioned image]), comprising 24.1k responses (one human-written and nine LM-generated per question) with native human preference using a 10-point scale on 2580 multilingual questions focusing on Chinese and Arab cultures (825 and 697 in respective languages; additional 1,058 covering other cultures). Compared with existing datasets Yin et al. (2022); Palta and Rudinger (2023); Shi et al. (2024); Chiu et al. (2024b), our data offers more balanced coverage and authentic viewpoints, especially for topics such as social norms and cultural commonsense, where previous works have limited data coverage. To achieve this, we have native speakers both re-annotate existing data resources, and curate new data on those rarely documented yet deeply embedded aspects of daily life.
Using CARE, we demonstrate that multilingual human cultural preference enhances LMs’ cultural awareness through alignment strategies, such as direct preference optimization (DPO, Rafailov et al., 2024), without compromising general capabilities. This cultural alignment proves beneficial beyond standard human preference datasets, with consistent gains across both chain-of-thought (CoT) and role-play prompting settings. We further analyze post-alignment performance across different LMs and find that for cultural alignment to be effective, base LMs need a foundational level of cultural knowledge. That is, cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs. We then examine the impact of cross-cultural alignment, and notice that aligning models with data from multiple cultures results in greater improvements than using single-culture data. This highlights the benefits of expanding geographical coverage in LM alignment.
Moreover, we evaluate the cultural awareness of large LMs (e.g. GPT-4o) in comparison to native speakers and web content. We observe cultural awareness gaps among LMs. For example, the more Chinese-centric Qwen2.5-72B excels on Chinese cultural entities, literacy, and social norms, while GPT-4o and Mistral-Large match or outperform it on opinion and commonsense about Chinese culture, likely due to differences in their training corpora. To contextualize these gaps, we analyze web content across languages. For questions on social norms and cultural commonsense, where native humans exhibit high confidence, Google search in native languages often return fewer relevant answers than in English. This indicates that such knowledge is often unstated in native contexts while more explicitly articulated in foreign languages Yin et al. (2022).
In summary, our key contributions include:
  • We introduce CARE, a multilingual multicultural human preference dataset focusing on Chinese and Arab cultures (§3).
  • We show that cultural alignment with CARE improves LMs across model families (§4).
  • Motivated by LM performance gaps, we analyze web content across languages and reveal documentation disparities (§5).
Figure 2: Example questions in Arabic and Chinese (with English translations) from the 5 cultural context categories in CARE Refer to caption: Cultural Entities, Cultural Opinion, Social Norms, Cultural Commonsense, and Literacy3.1).
2 Related Work
The widespread use of LMs has sparked research interest in their relevance to diverse cultures Adilazuarda et al. (2024); Shen et al. (2024); Pawar et al. (2024); Liu et al. (2024). Several studies investigate the cultural alignment of LMs by examining their responses to questions from social surveys that reflect human values and attitudes Haerpfer et al. (2021); Cao et al. (2023). It has been consistently shown that LMs favor answers associated with Western culture AlKhamissi et al. (2024); Abdulhai et al. (2023), even when prompted in different languages Masoud et al. (2023); Wang et al. (2023) or after undergoing preference optimization Ryan et al. (2024). While these survey-based studies assess general beliefs on broad cultural topics such as democracy or family values, they have limited coverage of detailed cultural nuances and phenomena such as local customs and daily practices.
Another line of work has developed culture-specific evaluation resources such as knowledge bases of cultural facts Keleg and Magdy (2023); Yin et al. (2022); Zhou et al. (2024b), entity-centric cultural benchmarks Naous et al. (2024), and user self-reported cultural experiences Shi et al. (2024). Other works have also constructed QA datasets for different cultural aspects such as culinary customs Palta and Rudinger (2023), norms Rao et al. (2024); Zhan et al. (2024), social etiquette Chiu et al. (2024a, b), and more Arora et al. (2024); Mousi et al. (2024). However, these are mainly designed for evaluating LMs, often in multi-choice QA or text-infilling formats, and are not readily suitable for improving LMs via human preference learning, which requires free-text QA and human ratings on model outputs.
Several multilingual instruction-tuning Muennighoff et al. (2022); Singh et al. (2024); Alyafeai et al. (2024) and preference resources Lian et al. (2023); Bartolome et al. (2023); Ahmadian et al. (2024) have been developed to improve multilingual performance. While they include culturally relevant samples, they primarily offer general instructions and universal preferences for safety and helpfulness. Few works Li et al. (2024a, b); Kirk et al. (2024) attempt to tune LMs on culturally relevant instructions, but they rely on synthesizing data with existing models and mainly express knowledge in English (more on this in §3.1). This may exacerbate biases in the pre-training data and miss authentic viewpoints. Unlike previous studies, our work focuses on culturally aligning LMs by learning from native speakers’ cultural preferences.
3 Constructing CARE [Uncaptioned image]
To study preference optimization of LMs for cultural awareness, we introduce CARE, which consists of 2,580 culture-specific questions and 24.1k responses (one human-written and nine LM-generated per question) with human ratings on a 10-point Likert scale. We begin by identifying and addressing the limitations in existing data resources (§3.1), and then further collect new data for cultural alignment (§3.2).
3.1 Limitations of Existing Data Resources
We start with the Aya dataset Singh et al. (2024), the largest multilingual instruction-tuning resource that contains human-written questions and answers. Though its natural instances collected from fluent native speakers offer authentic viewpoints, only part of its contents focus on culture, and the data quality varies for different languages. After manual inspection and filtering out generic questions (e.g. “How many hearts does an octopus have?”), we yield 1,324 (out of 4,909) question and answer pairs that are culturally relevant in Chinese but only 457 (out of 14,210) samples in Arabic. To expand the Arabic set, we apply the same filtering process to around 2,000 samples from CIDAR Alyafeai et al. (2024), a human-written Arabic instruction dataset, resulting in 162 relevant samples.
Figure 3: Overall coverage of our CARE benchmark per cultural category. CARE provides 16.6$\times$ more coverage of social norm and cultural commonsense questions compared to two existing instruction tuning datasets and four cultural knowledge bases (§3.1).
[…]
Moreover, our analysis reveals that existing resources provide highly unbalanced coverage of various cultural aspects, with a notable lack of data on social norms and commonsense, as shown in Figure 3. We manually classify all questions into one of five categories: (1) Cultural Entities, where the question asks about culture-specific entities Naous and Xu (2025), (2) Cultural Opinion, where the question asks about a subjective interpretation for a cultural entity, (3) Social Norms, where the question is about social interactions between individuals Huang and Yang (2023), (4) Cultural Commonsense, where the question is about daily phenomena that locals may take for granted Shen et al. (2024), and (5) Literacy, where the question is about language usage, proverbs, or slang Wuraola et al. (2024). Figure 2 shows examples for these cultural aspects. Inter-annotator agreement is 0.90, measured by Cohen’s Kappa, for this classification.
Finally, aligning LMs using preference optimization algorithms requires human judgments on both suitable and unsuitable responses, beyond only the existing ground-truth demonstrations.
3.2 Multilingual Cultural Alignment Data
To address the aforementioned shortcomings, we collect new non-English data for cultural alignment as follows:
Social Norm and Commonsense.
To expand the samples on social norms and cultural commonsense, we ask native Chinese and Arabic speakers who are international college students undergoing culture transfer experiences themselves to curate such samples manually. To help brainstorm, the speakers are instructed to leverage international as well as regional social media platforms or forums (e.g., Twitter, Reddit, Zhihu, RedNote), and search for posts where users describe their culture shift experiences using search keywords such as “most surprised abroad, culture shock, first time to the US, etc.”. Taking inspiration from such discussions, the speakers create 190 samples in Chinese and 196 samples in Arabic. This human-curated data is more authentic than synthetic data, preventing possible inaccuracy and overgeneralization (e.g. “In China, it is common to drink soup after the main dish.”) we have observed in the existing datasets.
Culture-specific Human Preference.
Given these culture-specific questions, gathering human preferences on the cultural relevance of responses is crucial to performing cultural alignment of LMs through preference optimization. To obtain these cultural preferences, we present the native annotators with the zero-shot responses of 9 different LMs for each question within CARE. We specifically use the instruct version of recent multilingual LMs of Llama3.3-70B Dubey et al. (2024), Qwen2.5-72B Yang et al. (2024), Gemma2-27B Team et al. (2024), Mistral-Large, and GPT-4o. We also use their smaller-sized Llama3.1-8B, Qwen2.5-7B, Gemma2-9B, and Mistral-7B.
We then instruct the annotators to rank the generated responses for each question from the most to least culturally appropriate and assign each response a rating on a scale of 1 to 10 (1: poor $\rightarrow$ 10: excellent). Those ratings reflect how well responses match the cultural expectations of native speakers, which we use to construct preference pairs for cultural alignment in our experiments (§4.1). For responses that deviate from answering the questions (e.g., generating similar questions or repetitive content), annotators do not assign a rating.
Human Annotation.
We hire two native Chinese speakers and two native Arabic speakers who are college students to collect and annotate all the data in CARE, as described in §3.1 and §3.2. Besides the 5-class cultural categories, we also annotate each sample with the Associated Culture for experiments in §4.5: Native (questions about the native culture; i.e. Chinese or Arab), Foreign (questions about other cultures), or General (questions that are not specific to a particular culture). For further fine-grained regional cultural knowledge evaluation (Appendix D.1), we also annotate samples for their Geographic Scope (sub-nationwide, nationwide, continent-wide, or worldwide). See Appendix A.1 for the annotation guideline.
Data Split.
For the experiments in §4 and §5, we construct culture-specific test sets with 150 questions each for Chinese and Arab cultures, randomly sampling 30 questions from each of the five cultural categories. The remaining 1,513 Chinese and 644 Arabic questions, along with their responses in CARE, are used to form the training set. This training data also includes questions about other foreign cultures beyond Chinese and Arab.
Chinese Arabic Model Entities Opinion Norms C. sense Literacy Average Entities Opinion Norms C. sense Literacy Average [Uncaptioned image] Gemma2-9B Vanilla 4.89 8.46 7.55 6.67 4.72 6.49 4.73 6.56 6.33 5.60 3.28 5.33 Aligned[Uncaptioned image] 5.50 8.53 7.90 7.26 5.17 6.89 5.30 6.82 6.93 6.55 3.57 5.84 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Llama3.1-8B Vanilla 3.14 4.16 4.62 3.93 3.03 3.78 4.08 3.62 2.87 3.07 2.07 3.30 Aligned[Uncaptioned image] 3.86 5.90 5.36 5.56 3.69 4.88 4.33 4.43 3.70 4.50 2.36 3.86 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Qwen2.5-7B Vanilla 6.89 7.86 7.48 6.80 7.37 7.28 4.65 5.84 5.44 4.88 2.84 4.61 Aligned[Uncaptioned image] 7.20 8.76 7.66 6.90 7.53 7.61 4.55 6.40 5.55 5.33 3.35 5.06 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Mistral-7B Vanilla 3.03 3.83 3.93 4.38 2.43 3.53 2.56 2.46 2.03 2.13 1.34 2.11 Aligned[Uncaptioned image] 2.43 3.90 4.53 5.00 2.20 3.61 2.60 3.36 2.46 2.10 1.40 2.38 [Uncaptioned image] [Uncaptioned image]
Table 1: Average scores (1: poor $\rightarrow$ 10: excellent) in responding to questions related to Chinese culture in Chinese and Arab culture in Arabic. Performances are presented for vanilla LMs and LMs after cultural alignment using DPO on CARE[Uncaptioned image]. For each LM, the row labeled “Vanilla” corresponds to the original model (gray plot), and “Aligned” is after cultural alignment (colored plot). “C. sense” stands for “Commonsense”.
[…]
Table 2: Performance comparison w.r.t. different preference learning algorithms on CARE[Uncaptioned image] data. Results show the average score improvements over the vanilla model.
4 Aligning LMs for Cultural Awareness
Enabled by CARE, we investigate whether a small amount of high-quality cultural alignment data can improve existing models through preference optimization. We first conduct cultural alignment on medium-sized LMs and show how much improvement we can achieve with CARE in §4.1 and §4.2. Next, we examine the impact of scaling data and model size (§4.3) and whether cultural alignment sacrifices LMs’ capabilities on general tasks (§4.4). Finally, we analyze the importance of including data from diverse cultures during alignment (§4.5).
4.1 Experiment Setup
We perform human preference learning for cultural alignment of medium-sized LMs (7$\sim$9B parameters) using CARE on up to 8 Nvidia A40 GPUs.
Human Preference Optimization.
We experiment with different preference learning algorithms, including DPO Rafailov et al. (2024), KTO Ethayarajh et al. (2024), and SimPO Meng et al. (2024). These algorithms require preference pairs that contrast two responses for each question: one culturally appropriate and one that is less appropriate. We use the highest- and lowest-scored responses from the human preference data collected in CARE. We find that all three offline preference algorithms perform similarly well, as shown in Table 3.2; subsequently, we present most of the experiment results in this paper with DPO unless otherwise specified.
Baselines.
First, we analyze the effectiveness of learning from cultural preference compared to general preference. We align LMs using existing human preference datasets, such as OpenOrca Lian et al. (2023) and UltraChat Bartolome et al. (2023), which focus on truthfulness, honesty, and helpfulness. For a fair comparison, we randomly sample an equal number of training pairs from these datasets as in CARE’s training set. Second, we examine the effectiveness of SFT on CARE versus Alpaca Cui et al. (2023); Chen et al. (2023), a general instruction-tuning dataset with Chinese and Arabic adaptations. Third, for prompting-based approaches, we explore CoT prompting Wei et al. (2022) and role-play prompting Kong et al. (2023). See prompts used in Appendix B.
LM-as-a-judge Evaluation and its Reliability.
For evaluation, we adopt the LM-as-a-judge Zheng et al. (2023) strategy where we prompt a powerful LM to evaluate how aligned the generated response is to the human reference response. We use GPT-4o as the LM judge and prompt it to provide reasoning and a score on a scale of 1 to 10. We validate this setup by comparing it to the ratings of the native-speaking human annotators, where we achieve a high Pearson correlation of 0.93 between human and GPT-4o ratings. More details and our evaluation prompts can be found in Appendix C.
[…]
Table 3: Average scores (1: poor $\rightarrow$ 10: excellent) on Chinese and Arab cultures for a variety of prompting approaches, fine-tuning, and alignment using cultural (CARE) vs. generic instruction-tuning (Multilingual Alpaca) and preference (OpenOrca/UltraChat) data. SFT is performed on the instruction data only, while alignment conducts DPO on the human preference pairs. Aligning with cultural preferences enhances cultural awareness more effectively than using general preferences, with this improvement remaining consistent across various prompting setups.
4.2 Main Results
Table 1 presents the results of medium-sized LMs before and after cultural alignment using CARE, while Table 3 compares these results with various baselines. We have the following key findings:
Using CARE, culturally-aligned LMs achieve higher average scores (up to 29% improvement) in both Chinese and Arabic compared to the vanilla checkpoints, even for these already instruction-tuned models. This resonates with the findings of Zhou et al. (2024a) and shows that a relatively small amount of carefully curated data by humans can improve LMs’ alignment. We also see a noticeable gap among LMs developed by different regions. The Qwen2.5-7B developed by China-based Alibaba performs the best in Chinese and can be further improved from 7.28 to 7.61 (out of 10) overall by aligning with CARE. Interestingly, the aligned version of Gemma2-9B outperforms Qwen2.5-7B on social norms and commonsense for Chinese (more on this in §5.2). We also test its generalization capabilities on other culture-related tasks in Appendix D.
Cultural alignment strengthens stronger LMs but fails to address weaknesses in weaker LMs. Qwen2.5-7B and Gemma2-9B, the top-performing LMs on Chinese and Arabic, respectively, show consistent improvement across all five cultural categories after alignment. However, alignment shows limited gains where the model’s initial performance is poor. Mistral-7B is not improved on Chinese entities and literacy, where its starting scores are only 3.03 and 2.43. Similarly, Qwen2.5-7B, does not benefit from alignment on Arabic entities. This suggests that base models need a foundational level of cultural knowledge for alignment to be effective.
Culture-specific human preference data is helpful beyond general human preference. Table 3 shows that alignment with CARE’s cultural preference consistently outperforms general preference datasets by up to 42%. Additionally, cultural alignment with CARE brings improvement in both CoT and role-play prompting. While culture-specific SFT outperforms generic instruction tuning, alignment yields greater gains, highlighting the importance of human cultural preference data over ground-truth demonstrations alone.
4.3 How does scaling affect alignment?
We examine how scaling both model size and data volume influences the performance before and after cultural alignment. Specifically, we consider the Qwen2.5 {1.5B, 3B, 7B, 14B} and the Gemma2 {2B, 9B} series, which offer multiple model sizes. For data sizes, we align models using DPO on different proportions {0%, 25%, 50%, 75%, 100%} of Chinese cultural preference pairs within CARE.
As shown in Figure 4, consistent improvements are observed across different model and data sizes. Even a relatively small amount (e.g., 25%) of data in CARE can lead to improvements, particularly for smaller-sized LMs. Scaling up the data leads to progressively better performance, highlighting the potential benefits of employing more cultural preference data in future work.
Figure 4: Impact of model size and preference data volume on cultural awareness performance. The average scores (1: poor $\rightarrow$ 10: excellent) of aligned models are plotted against different % of preference pairs in CARE. Improvements are achieved across different model sizes and data sizes, in comparison to the vanilla model (0%).
4.4 Does alignment impact general capability?
We then examine whether cultural alignment impacts the model’s overall knowledge and capabilities, using well-established benchmarks for Chinese, Arabic, and English: ArabicMMLU Koto et al. (2024), ChineseMMLU Li et al. (2023), MMLU Hendrycks et al. (2020), TruthfulQA Lin et al. (2021), and WinoGender Rudinger et al. (2018). We compare different versions of Llama3.1-8B-Instruct: the vanilla LM, the aligned LM with general human preference from the combined UltraChat and OpenOrca datasets (totaling 3,718 samples), the culturally aligned LM with cultural human preference from CARE, and the aligned LM with both general and CARE’s cultural preference.
The results in Figure 5 show very small differences between the vanilla LM and its aligned version, on benchmarks in both native language and English, suggesting that cultural alignment does not hinder the model’s overall capabilities.
Figure 5: Accuracy of Llama3.1-8B-Instruct variants on general benchmarks assessing overall capabilities in native languages (Chinese and Arabic) and English. Slight differences are observed among the LM variants after applying different alignment setups.
4.5 Should we align with only native data?
We analyze the impact of incorporating data about different cultures in the alignment process. To ensure a fair comparison, we select an equal number of samples for each trial. Specifically, we align Llama3.1-8B-Instruct with 676 Chinese and 546 Arabic pairs sampled from three different contexts in CARE: native (Chinese or Arab) culture, foreign cultures (non-Chinese or non-Arab), and mixed pairs (half native and half foreign).
Results are shown in Figure 6. We find that performance drops on native culture when the model is aligned using only foreign culture samples, but improves when aligned with the same native culture samples (4.69 $\rightarrow$ 4.87 in Chinese and 2.98 $\rightarrow$ 3.50 in Arabic); interestingly, performance is further improved when the model is aligned by combining native and foreign cultural data. This indicates that expanding geographical coverage in alignment could further enhance cultural awareness.
[…]
Table 4: Average scores on CARE of larger LMs and web content, evaluated by the judge LM across all samples (1: poor $\rightarrow$ 10: excellent). “Human Awareness (%)” indicates the percentage of questions for which natives know the correct answer. Qwen2.5 developed by Chinese-based Alibaba excels on Chinese entities, norms, and literacy, while GPT-4o and Mistral-Large are comparably good at opinions and commonsense. For social norms and cultural commonsense, where native humans show high confidence, Google search often returns fewer relevant answers.
5 Prompting LMs for Cultural Awareness
Figure 6: Average scores of Llama3.1-8B-Instruct on native and foreign cultures when aligned using data from native, foreign, or mixed cultures. The highest performance on native culture is achieved when mixing native and foreign samples during alignment.
Beyond improving medium-sized LMs with culture-specific human preference, we evaluate the performance of larger LMs and compare with native human familiarity and web-based sources (§5.1). To better understand these differences, we examine the role of querying language (§5.2).
5.1 Main Results
We follow the evaluation setup in §4.1 and assess the larger ($>$ 25B parameters) versions of the evaluated LMs with zero-shot prompting. Table 4 shows the results. Larger models are better for both Chinese and Arabic, with GPT-4o achieving the best overall performance. The more Chinese-centric model Qwen2.5-72B excels in the knowledge about cultural entities (8.42), social norms (9.06), and literacy (8.53) for Chinese; meanwhile, GPT-4o and Mistral-Large are comparably good if not better at offering opinions (9.43 and 8.96) and answering commonsense questions (9.00 and 8.58) about Chinese culture. This is interesting and likely because cultural commonsense knowledge is often unstated in the native language, while more balanced and thoughtful opinions may be expressed in foreign languages. To illustrate and test this hypothesis, we conduct further analysis in §5.2.
Human Awareness.
To assess human familiarity with the culture-specific questions, we ask native-speaking annotators whether they knew the correct answers without looking them up when presented with questions from CARE. Table 4 shows the percentages of questions humans could answer immediately with confidence, which are very high for social norms and commonsense knowledge.
Search Engine.
To further examine the coverage of cultural knowledge in real-world text, we extract the main content from the top-10 ranked webpages returned by Google [^1] Google programmable search engine API: https://programmablesearchengine.google.comand evaluate their quality using the same LM-as-a-judge assessment framework, recording the highest score for each question in the CARE. Results are shown in Table 4. Since retrieved sentences from relevant web pages often fail to answer the questions directly, their scores are generally lower than LMs’ scores.
(a) Average scores achieved by Qwen2.5-72B-Instruct and GPT-4o when prompted in native languages versus English.
(b) Average scores of retrieved webpage main content when searching CARE questions in native languages versus English.
Figure 7: Performance comparison of cultural awareness in different languages. Comparison between (a) LM performance and (b) retrieved webpage content, showing that native-language questions capture cultural nuances for literacy, opinion, and entities. However, the gaps narrow or sometimes reverse for cultural norms and commonsense.
5.2 Cross-lingual Analysis
We translate all questions and responses from CARE that are written in Chinese and Arabic into English to enable cross-lingual evaluation.
Figure 7(a) shows that LMs generally perform better on culture-specific questions when prompted in the native languages than in English, but more notably for Chinese than for Arabic. This advantage is also especially obvious for the literacy category, while the gap narrows in categories such as cultural commonsense. A possible explanation is that English corpora contain fewer discussions about Chinese or Arab cultures.
We also examine how language choice impacts search engine results. As shown in Figure 7(b), cultural information on entities, opinions, and literacy tends to be of higher quality when retrieved in native languages. Interestingly, this trend reverses for questions related to cultural commonsense and social norms, such as “In China, does the ticket time indicate when the feature movie starts, or are there trailers played before the main feature?”, where English searches yield higher-quality content. This aligns with the potential documentation gap mentioned in §5.1, where native speakers often assume everyday cultural knowledge is implicit, whereas non-native speakers are more likely to actively seek out and document it. A similar pattern is observed for Arabic. For a small number of Arabic questions (10% opinion, 14% social norms, 14% commonsense, and 7% literacy), the search engine also failed to return any results. One such example is “A Jordanian friend mentioned that it’s polite not to open a gift in front of the giver. Is this true?”.
6 Conclusion
We introduce CARE, a human-written, multilingual resource comprising 2,580 culture-specific questions about Chinese and Arab cultures, along with responses that include human preferences. Through extensive experiments, we investigate that a small amount of high-quality cultural alignment data can improve existing LMs via preference optimization, across different model families and scales. This cultural-specific alignment offers benefits that go beyond generic human preference and proves effective under various prompting setups. Our analyses also reveal gaps in cultural awareness among LMs, native-speaking humans, and search engines. For example, cultural commonsense is often left implicit in the native language but is more explicitly mentioned in foreign-language content. By releasing CARE, we hope to encourage the development of more inclusive and culturally adaptive LMs.
Limitations
In this work, as we investigate how to improve LMs by aligning with multilingual multicultural human preference data, we also identified a couple of areas for further exploration. First, while filtering the Arabic portion of the Aya dataset, we notice a large number (3,831) of questions centered on Islamic knowledge and religious history. We exclude them, as religion-specific understanding in LMs falls outside the scope of our study that focuses on culture-specific phenomena and factual knowledge, as defined in §3.1. Nonetheless, future research on religious QA or alignment could adopt a similar approach to CARE by collecting preference data from credible religious sources.
Second, LMs may occasionally respond to culture-specific questions in a stereotypical or biased manner, stemming from a lack of cultural understanding. For instance, when asked “In China, what does the term “laowai” mean when referring to foreigners?”, LMs often incorrectly interpret the term as disrespectful. In reality, “laowai” is typically a neutral descriptor in Chinese, used to denote foreigners without negative connotations. While CARE includes examples that clarify such misunderstandings through human-annotated responses, future studies in cultural red-teaming can further investigate and address these failure modes.
Acknowledgments
The authors would like to thank Nour Allah El Senary, Jad Matthew Bardawil, Chao Jiang, and Yuming Pan for their help in data annotation; Yao Dou and Xiaofeng Wu for their helpful feedback. We would also like to thank Microsoft’s Azure Accelerate Foundation Models Research Program and NVIDIA’s Academic Grant Program for providing computational resources to support this work. This work was funded in part through a Sony Faculty Innovation Award, and by the NSF under grant number IIS-2144493, IIS-2052498 and SMA-2418946. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A CARE: Details
A.1 Annotation Details
To construct the question-answer pairs in CARE, we recruited two native Chinese speakers and two native Arabic speakers who are undergraduate students. The annotators were paid at the rate of 16 USD/hour and told that they were helping create and annotate data for improving the cultural awareness of LMs. The detailed annotation guideline for dataset construction is provided in Figures 16 to 20, and the instruction for preference annotation is provided in Figures 14. We also show our rank-and-rate web interface for collecting human cultural preference pairs in Figure 15.
A.2 Statistics
Table 5 shows the detailed statistics for Chinese and Arab culture in CARE across the cultural categories, stratified by the data source. The questions specific to Chinese culture are written in Chinese, while the questions specific to Arab culture are written in Arabic. The culturally-relevant samples obtained from the instruction-tuning dataset mostly fall within the cultural entities, cultural opinion, or literacy categories, while samples obtained from cultural knowledge bases provide more social norms and cultural commonsense data.
Table 6 shows the detailed statistics for other foreign cultures in CARE for each cultural category. These samples were obtained when filtering the instruction tuning datasets and cultural knowledge bases (§3.1). We wrote these foreign samples in both Chinese and Arabic versions and used them as part of our training set and in our analysis on the impact of the source culture in alignment (§4.5). We also manually collected more samples for social norms and cultural commonsense from online websites to cross 100 samples in each category. We ensured these foreign samples do not include anything relevant to Chinese or Arab cultures.
Table 7 shows the detailed statistics for Chinese and Arabic samples in CARE across the geographic scope. Most samples fall into nationwide category, while sub-nationwide samples provide insights about more detailed cultural information, continent-wide and worldwide samples assess understanding between several cultures.
Appendix B Implementation Details
DPO Alignment.
We perform preference optimization using DPO for 3 epochs until the loss converges. Training is done with a batch size of 128 on 4-8 NVIDIA A40 GPUs. We tune the learning rate in the range $\{1e^{-7},5e^{-7},9e^{-7}\}$ and set beta for DPO as 0.1. The training involves full fine-tuning with 5 warmup steps and employs a linear learning rate scheduler.
LM Inference.
For open-sourced LMs, we run inference on 1 NVIDIA A40 GPU with the vLLM library [^2] https://docs.vllm.ai Kwon et al. (2023). We perform decoding by setting the following parameters {temperature=0.7, top_p=1}. We limit the context length by setting {max_model_len=2048}. We also limit the number of generated tokens by the models by setting {max_tokens=1024}. For the closed-source GPT-4o LM, we run inference with Azure OpenAI API.
Baselines Prompt Templates.
The prompt templates used for our prompting-based baselines (§4.2) are provided in Figure 8 (role-play prompting) and Figure 9 (CoT prompting).
Role-play Prompt Template
You are a native [Chinese/Arab] person, familiar with
[Chinese/Arab] culture and traditions.
{Question}
Figure 8: Prompt template for role-play inference. The LM is told to take on the persona of a native [Chinese/Arab] person who is familiar with the culture’s traditions, and then asked the culture-specific {Question}.
CoT Prompt Template
You are a helpful assistant.
{Question}
Let’s think step by step.
Figure 9: Prompt template for CoT inference. We provide the LM with the test question in {Question}, then ask it to think step by step when providing the answer.
Chinese Arabic Model Sub-nationwide Nationwide Continent-wide Worldwide Sub-nationwide Nationwide Continent-wide Worldwide [Uncaptioned image] Gemma2-27B 6.50 7.16 8.19 6.47 4.71 6.73 4.88 6.29 [Uncaptioned image] Llama3.3-70B 6.20 6.92 7.68 5.57 5.14 6.16 4.31 5.50 [Uncaptioned image] Qwen2.5-72B 7.70 8.49 8.60 6.78 7.86 7.62 5.19 8.58 [Uncaptioned image] Mistral-Large 7.42 8.11 8.73 6.70 5.29 7.24 5.31 8.08 [Uncaptioned image] GPT-4o 8.59 8.74 8.72 7.11 7.43 8.15 7.79 8.92
Table 8: Performance comparison w.r.t. geographic scope. Scores are computed on the entire CARE data, including both native culture and foreign culture.
Appendix C LM-as-a-Judge
C.1 Evaluation Prompts
We instruct GPT-4o as the judge LM to score a model’s response to culture-specific questions in CARE. For each cultural category, we provide the judge LM with a detailed evaluation guideline, the culture-specific question, the generated response, and the human reference response, and ask it to score the response on the 1-10 scale. Our evaluation prompt templates are provided in Figure 11 (Entities & Opinion), Figure 12 (Norms & Commonsense), and Figure 13 (Literacy).
C.2 Correlation with Human Ratings
Fig 10 shows the rating distribution of GPT-4o and native speakers. We can see a clear correlation between both rating distributions. We also calculate correlation metrics and obtain a Pearson correlation of 0.933, Spearman correlation of 0.901, and Kendall’s Tau correlation of 0.733, all indicating high agreement between the judge LM ratings and human ratings. One difference we notice is that humans tend not to assign extreme ratings (1 or 10), yet the LM judge has a higher frequency of those ratings. This is mostly noticeable at the 9-10 rating range where the LM ratings are more equally distributed between a rating of 9 and 10, but the human ratings mostly consist of 9.
Figure 10: Rating score distributions of the judge LM (GPT-4o) and native human evaluators. The judge LM highly correlates with human ratings.
Appendix D Additional Results
D.1 Performance by Geographic Scope
Table 8 shows the results achieved by different LMs when stratified by geographic scope. We find that the models struggle the most with sub-nationwide questions. Most models achieve the best scores on continent-wide or worldwide questions.
D.2 Performance on other cultural tasks
We also evaluate the performance of LMs after cultural alignment with CARE on other cultural tasks using out-of-domain culture-related data. Specifically, we use the Include benchmark Romanou et al. (2025), as it offers multilingual versions for each culture and employs a multi-choice format, enabling us to assess the robustness of our alignment. From Table 9, we observe that alignment with culture-specific human preference can also benefit the performance on the Include benchmark, indicating the generalization capability of CARE.
[Uncaptioned image] Gemma2-9B [Uncaptioned image] Llama3.1-8B [Uncaptioned image] Qwen2.5-7B [Uncaptioned image] Mistral-7B Vanilla CARE Vanilla CARE Vanilla CARE Vanilla CARE Include (zh) 57.79 59.44 51.00 54.68 75.22 75.59 40.36 40.73 Include (ar) 57.42 54.71 50.54 51.08 56.15 55.97 35.32 36.95
Table 9: Accuracy (%) on the Include benchmark for Chinese and Arabic subsets, before and after cultural alignment with CARE. Improvements in the Include benchmark can also be observed after alignment with CARE.
Appendix E Responsible NLP Checklist
In our work, we ensure that the use of existing artifacts is consistent with their intended use. Specifically, all samples collected for the CARE dataset are sourced from open-sourced datasets, adhering strictly to the original access conditions intended for research purposes. We meticulously filter the data to remove any instances of toxic or offensive language and ensure that it does not contain any personally identifiable information. For the artifacts we create, we clearly specify their intended use within research contexts, maintaining compatibility with the original access conditions. We acknowledge the use of GPT-4 for grammar checking and word polishing.
Cultural entities and opinion evaluation
Please serve as an unbiased evaluator and critically assess the quality of the assistant’s response to the user’s question
presented below. When evaluating, focus on the following aspects:
1. **Accuracy**: Is information in the response factually correct and up-to-date?
2. **Depth**: Does the response demonstrate a deep understanding of the topic, providing thorough explanations and context?
3. **Relevance**: Does the response stay focused on the specific question without including unnecessary information?
Begin your evaluation with a brief commentary explaining your judgment on each of these criteria. Aim to be as objective as
possible in your assessment.
After providing your commentary, assign a numerical rating to the response on a scale from 1 to 10, where:
- **1-2**: Poor — The response is highly inaccurate, lacks detail, contains significant incorrect information, and/or
includes irrelevant information.
- **3-4**: Below Average — The response is partially accurate, addresses some parts of the question but lacks detail, and
may include irrelevant information.
- **5-6**: Average — The response is moderately accurate but may contain minor errors, addresses most parts of question with
adequate detail, and is mostly relevant.
- **7-8**: Good — The response is mostly accurate, addresses all parts of the question with good detail, and is relevant
with minimal irrelevant information.
- **9-10**: Excellent — The response is highly accurate, provides comprehensive detail, and contains no irrelevant information.
Please format your rating as follows: "Rating: [[number]]". For example: "Rating: [[6]]".
## Question: {question}
## Golden answer: {answer}
## Assistant’s response: {response}
Figure 11: LM-as-a-judge prompt template for cultural entities and opinion questions.
Social norms and commonsense evaluation
Please serve as a critical evaluator and rigorously assess the quality of the assistant’s response to the user’s question
shown below. When evaluating, prioritize the following stringent criteria:
1. **Accuracy**: Based on the golden answer, is the information factually correct and representative of real-world cultural
context?
2. **Comprehensiveness**: Does the response provide clear explanations and consider possible nuances or exceptions within
the cultural context?
3. **Relevance**: Is the response perfectly tailored to the specific cultural context, without any generalizations or
inaccuracies?
Begin your evaluation with a detailed commentary critically analyzing each of these criteria. Strive to be as objective
and discerning as possible in your assessment.
After providing your commentary, assign a numerical rating to the response on a scale from 1 to 10, where:
- **1-2**: Poor — The response fails to meet basic expectations for accuracy or relevance, showing major misunderstandings or
errors.
- **3-4**: Below Average — The response has substantial inaccuracies or omissions, only partially addressing the user’s needs.
- **5-6**: Average — The response is fairly accurate and relevant but lacks depth, missing important details or subtleties.
- **7-8**: Good — The response is accurate and covers most aspects well, though it may lack in minor details or perfect
contextual alignment.
- **9-10**: Excellent — The response is outstanding in all respects; it is precise, detailed, fully relevant, and excellently
contextualized.
Please format your rating as follows: "Rating: [[number]]". For example: "Rating: [[6]]".
## Question: {question}
## Golden Answer: {answer}
## Assistant’s response: {response}
Figure 12: LM-as-a-judge prompt template for social norms and commonsense questions.
Literacy evaluation
Please serve as a critical evaluator and rigorously assess the quality of the assistant’s response to the user’s question shown
below. When evaluating, prioritize the following stringent criteria:
1. **Accuracy**: Is the information in the response factually correct and contextually appropriate?
2. **Interpretation**: Does the response offer insightful and well-supported interpretations of the literary work or topic?
3. **Textual Evidence**: Does the response appropriately reference and analyze specific parts of the text to support its points
when necessary?
4. **Relevance**: Does the response stay focused on specific question without including unnecessary information?
Begin your evaluation with a detailed commentary critically analyzing each of these criteria. Strive to be as objective and
discerning as possible in your assessment.
After providing your commentary, assign a numerical rating to the response on a scale from 1 to 10, where:
- **1-2**: Poor — The response fails to meet basic expectations for accuracy or relevance, showing major misunderstandings or
errors.
- **3-4**: Below Average — The response has substantial inaccuracies or omissions, only partially addressing the user’s needs.
- **5-6**: Average — The response is fairly accurate and relevant but lacks depth, missing important details or subtleties.
- **7-8**: Good — The response is accurate and covers most aspects well, though it may lack in minor details or perfect contextual
alignment.
- **9-10**: Excellent — The response is outstanding in all respects; it is precise, detailed, fully relevant, and excellently
contextualized.
Please format your rating as follows: "Rating: [[number]]". For example: "Rating: [[6]]".
## Question: {question}
## Reference Answer: {answer}
## Assistant’s response: {response}
Figure 13: LM-as-a-judge prompt template for literacy questions.
Figure 14: Cultural preference annotation instructions.
Figure 15: The interface for annotating culture-specific human preference.
Figure 16: Annotation Guideline (1/5).
Figure 17: Annotation Guideline (2/5).
Figure 18: Annotation Guideline (3/5).
Figure 19: Annotation Guideline (4/5).
Figure 20: Annotation Guideline (5/5).